OpenStudy (anonymous):

HElp please Medal Rectangle A has an area of 9 - x2. Rectangle B has an area of x2 + x - 12. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (anonymous):

(9-x^2) / (x^2 + x -12) (3 + x) (3 - x) / (x + 4)(x - 3) taking out the negative sign from (3 - x) gives - (3 + x)(x - 3) then cancel the ( x - 3 ) -(x+3)/(x+4) is the answer.

OpenStudy (anonymous):

Oh sure, if I know how to solve it :)

OpenStudy (anonymous):

120r^3 / sq rt (16r^9) = 120r^3 / 4r^4 sq rt r = 30 / r sq rt r rationalize the denominator 30 sqrt r/r^2

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

(3 + x) (3 - x) / (x + 4)(x - 3) (3 - x) = (-x + 3 ) = - ( x - 3 ) Therefore -( 3 + x) ( x - 3 ) / (x + 4)(x - 3)

OpenStudy (anonymous):

You may get reported by a moderator

OpenStudy (anonymous):

You are medal fishing and typing answers